OpenStudy (srtauser):

Solve. 3x2 = -15x + 21

OpenStudy (soprano.h.d0816):

Hey there!!! Since you are new here, read this tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8 You might not have known, but you are in the "OpenStudy Feedback" section. Place your mouse over where it says OpenStudy Feedback, and find the correct subject by clicking the 'Find More Subjects' button or simply clicking the correct subject in the list that you have already studied! Please enjoy being on OpenStudy. And don’t forget to read OpenStudys’ Code of Conduct and also, OpenStudy's Terms and Conditions :D This section is for providing your Feedback/Appreciation regarding the site and by Feedback, we mean "helpful information or criticism that is given to someone to say what can be done to improve a product( which is the site in this case)" You can use this section: ⇝ If you want to appreciate to OpenStudy ⇝ To share your thoughts on how to improve OpenStudy ⇝ To inform the site about any difficulty you are facing while using OpenStudy ⇝ To ask any other question related to this site. ⇝ If you want to post any posts such as congratulating someone, saying Happy Birthday, etc. If you have any questions, please feel free to PM me or tag me in your post :D

OpenStudy (srtauser):

So i know what to do up until the point where my answer looks like 3x^2-15x-21=0

OpenStudy (soprano.h.d0816):

You need to add 21 to both sides

OpenStudy (srtauser):

then what? is it supposed to look like 3x^2-15x=21?

OpenStudy (soprano.h.d0816):

Yes. Now, 3x^2 is 9x-15x=21 ^^^^^^ Combine -6x = 21 --- --- -6 -6 x = ??

OpenStudy (srtauser):

-3.5

OpenStudy (srtauser):

what is that in a square root form?

OpenStudy (soprano.h.d0816):

Errr.... idk right off

OpenStudy (srtauser):

well my answer options are A. the quantity of negative 5 plus or minus i square root 3 all over 2 B. the quantity of 5 plus or minus square root 53 all over 2 C. the quantity of 5 plus or minus 3 square root 3 all over 2 D. the quantity of negative 5 plus or minus square root 53 all over 2

OpenStudy (soprano.h.d0816):

Errrr.....

OpenStudy (srtauser):

im stuck on it too

OpenStudy (soprano.h.d0816):

@agent0smith ??

OpenStudy (agent0smith):

3x^2-15x-21=0 From here, you can take out a common factor of 3.

OpenStudy (srtauser):

ok so 3(x-5x-7)=0

OpenStudy (srtauser):

I dont know what to do nexst because this trinomial can't be changed

OpenStudy (agent0smith):

You forgot the x^2 3(x^2-5x-7)=0 Now divide both sides by 3. You can't factor it, use the quadratic formula.

OpenStudy (srtauser):

then I get 5+_ sqrt 109 over 6

OpenStudy (agent0smith):

Keep in mind that you are using the quadratic formula for x^2-5x-7=0... it looks like you didn't do it for that.

OpenStudy (srtauser):

How do I do it for that then?

OpenStudy (agent0smith):

By using the quadratic formula. It looks you used 3(x^2-5x-7)=0 and not x^2-5x-7=0... meaning you did not divide both sides by 3.

OpenStudy (srtauser):

oh ok -5/3x-7/3=0 is this right?

OpenStudy (agent0smith):

3(x^2-5x-7)=0 When you divide both sides by 3, you do not touch anything inside the parentheses.

OpenStudy (srtauser):

so it is like i said before x^2-5x-7=0 or not?

OpenStudy (agent0smith):

Yes. Now use the quadratic formula.

OpenStudy (srtauser):

which is what I used before?

OpenStudy (agent0smith):

Then how did you get the "over 6" part? The denominator of the quadratic formula is 2a.

OpenStudy (srtauser):

oh ok. I see so it is (5+-√53)/2 Thank you so much!

OpenStudy (soprano.h.d0816):

He's very smart(:

OpenStudy (agent0smith):

Much better. And ty!